OpenStudy (anonymous):

The lengths of the sides of a square are increasing at a constant rate of 2ft/min. In terms of the perimeter, P, what is the rate of change of the area of the square in square feet per minute?

OpenStudy (anonymous):

let L represent length, A represent Area P = 4L A = L^2 dA/dl = 2L dL/dt = 2ft/min dA/dt = dA/dl * dl/dt = 2L (2) = 4L ft/min = P ft/min

OpenStudy (anonymous):

thanks!